OpenStudy (anonymous):

For every problem answered correctly on an exam, 3 points are awarded. For every incorrect answer 4 points are deducted. In a 10-question test a student scored 16 points. How many correct and incorrect answers did he have on the exam?

OpenStudy (mertsj):

r+w=10 3r-4w=16

OpenStudy (anonymous):

ok did that got to x =30+12y

OpenStudy (anonymous):

let x = correct answers, and y = incorrect ansewrs we can make these equation from the question: x+y=10 which is the number of questions given and, 3x-4y=16 which is the amount of score you get and then solve them. you can use calculator to solve them out :D

OpenStudy (mertsj):

1. Multiply equation 1 by 4. 2. Add the result of step 1 to equation 2 3. Divide both sides of the answer to step 2 by 7

OpenStudy (anonymous):

youll get x=10-y, then sub into second equation: 3(10-y) - 4y = 16, 30 - 3y -4y = 16 7y = 14, y = 2 then sub the value back in eq 1, x + 2 = 10 x=8

OpenStudy (anonymous):

Thank you! You have no idea how long I was stuck on that! Bless you both! :)

OpenStudy (anonymous):

no problem :D